Capacitors of C1 = 9 micro-farads, C2 = 8 micro-farads, and C3 = 8 micro-farads are connected purely in parallel to a 9 volt battery. What is the stored energy in C2 in micro-joules?
Answer #1


IN parallel combination the potential v across the two capacitors is same

V1 = V2 = 9 V

energy store in capacitor = U = (1/2)*C*V^2


U2 = (1/2)*C2*V2^2

U2 = (1/2)*8*10^-6*9^2 = 324 micro-joules